PLANE FLIES POINT A TO POINT B, 120 MILES IN 12 HOURS, FLYING INTO A HEADWIND. ON THE RETURN TRIP WITH THE SAME WIND AS A TAIL WIND, THE TRIP FROM POINT B TO POINT A TAKES ONLY 2 HOURS. HOW FAST IS THE PLANE FLYING THROUGH THE AIR? HOW FAST IS THE WIND BLOWING?
Answer by mathmaven53(29)

Let v be speed of plane through air and w the speed of the wind
From A to B the plane is traveling at speed v-w with the headwind
It takes 12 hours to go 120 miles
12(v-w) =120
v-w = 10
From B to A the plane is traveling at speed v+w with the tail wind
It takes 2 hours to go 120 miles
2(v+w) = 120
v + w = 60
Solve the system v-w = 10, v + w = 60